Two seconds after firing a rifle at a target the shooter hears the impact of the bullet on the target. Sound travels at 1100 feet per second, and the bullet at 1865 feet per second. Determine the distance to the target (to the nearest foot)

Answer by TimothyLamb(4379)

d = distance from shooter to target
s = d/t
---
time of bullet traveling, from shooter fires to bullet hits target:
use speed of bullet:
a = d/s
a = d/1865
---
time of sound traveling, from bullet hits target to shooter hears bullet hit target:
use speed of sound:
b = d/s
b = d/1100
---
total time elapsed, from shooter fires to shooter hears bullet hit target:
a + b = 2
d/1865 + d/1100 = 2
1100d/(1100*1865) + 1865d/(1100*1865) = 2
1100d + 1865d = 2*(1100*1865)
(1100 + 1865)d = 2*(1100*1865)
d = 2*(1100*1865)/(1100 + 1865)
d = 1383.81112985
---
answer:
distance from shooter to target = 1384 feet
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---